OpenStudy (anonymous):

Solve the following system of equations: 2x + 2y + z = 10 3x − y + 3z = 10 2x + 3y − 2z = 6

OpenStudy (welshfella):

use the elimination method start off by multiply the 3rd equation by -1 and add it to the first equation this will eliminate x

OpenStudy (welshfella):

2x + 2y + z = 10 -2x- 3y+ 2z = -6 add y + 3z = 4

OpenStudy (anonymous):

all 3 of them?

OpenStudy (welshfella):

no - now get another equation in y and z using the second and third equations

OpenStudy (welshfella):

multiply second equation by -2 and third equation by 3 then add

OpenStudy (welshfella):

sorry gotta go

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

I still don't get it . . . .

OpenStudy (anonymous):

how do I use the elimination one

OpenStudy (texaschic101):

first take 2 of the 3 equations...any 2.. and eliminate a variable 2x + 2y + z = 10 --->(-1) 2x + 3y - 2z = 6 --------------- -2x - 2y - z = -10 (result of multiplying by -1) 2x + 3y - 2z = 6 ---------------add y - 3z = -4 are you still with me...do you have any questions so far ?

OpenStudy (anonymous):

im here

OpenStudy (anonymous):

once you find that number where do you put it? back in the x?

OpenStudy (texaschic101):

now we take 1 of the equations we used and he one we did not use, and eliminate the same variable 2x + 2y + z = 10 --->(3) 3x - y + 3z = 10 --->(-2) -------------- 6x + 6y + 3z = 30 (result of multiplying by 3) -6x + 2y - 6z = -20 (result of multiplying by -2) -------------add 8y - 3z = 10 still following me ? any questions so far ?

OpenStudy (texaschic101):

now we take our results from the equations we did..and eliminate a variable y - 3z = -4 --->(-1) 8y - 3z = 10 ----------- -y + 3z = 4 (result of multiplying by -1) 8y - 3z = 10 -------------add 7y = 14 -- divide both sides by 7 y = 2 still there ? any questions so far ?

OpenStudy (anonymous):

okay

OpenStudy (texaschic101):

now since we have found y, we sub it back into one of the equations we have already eliminated the x from. y - 3z = -4 2 - 3z = -4 -3z = -4 - 2 -3z = -6 -- divide both sides by -3 z = 2 understand so far ?

OpenStudy (texaschic101):

now that we have found y and z, we sub those back into one of the original equations to find x. 2x + 2y + z = 10 2x + 2(2) + 2 = 10 2x + 4 + 2 = 10 2x + 6 = 10 2x = 4 x = 2 but before you turn in these answers, lets check them 2x + 3y - 2z = 6 2(2) + 3(2) - 2(2) = 6 4 + 6 - 4 = 6 6 = 6 (correct) so x = 2, y = 2, and z = 2 It is not really that hard...just very time consuming. And if you mess up at the beginning of the problem, it follows you all the way to the end. That is why it is always best to check your answers.

OpenStudy (texaschic101):

ANY questions at all ?

OpenStudy (texaschic101):

look over this real well and if you have any questions, just tag me :)

OpenStudy (anonymous):

Thank you so much! it really helped me you deserve a medal :)